# A Wavefield Correlation Approach to Improve Sound Speed Estimation in Ultrasound Autofocusing

### v1

@@ Abstract
Aberration often degrades ultrasound image quality when beamforming does not account for wavefront distortions. In the past decade, local sound speed estimators have been developed for distributed aberration correction throughout a medium. Recently, iterative sound speed optimization approaches have achieved more accurate estimates than earlier approaches, but these newer methods still struggle with decreased accuracy for media with reverberation clutter and large sound speed changes. To address these challenges, we propose using a wavefield correlation (WFC) beamforming approach when performing sound speed optimization. WFC correlates simulated forward-propagated transmit wavefields and backwards-propagated receive wavefields in order to form images. This process more accurately models wave propagation in heterogeneous media and can decrease diffuse clutter due to its spatiotemporal matched filtering effect. This beamformer is implemented using auto-differentiation software to then perform gradient descent optimization, using a total-variation regularized common midpoint phase focus metric loss, on the local sound speed map used during beamforming. This approach is compared to using delay and sum (DAS) with straight-ray time delay calculations in the same sound speed optimization approach on a variety of simulated, phantom, and in vivo data with large sound speed changes and clutter. Results show that using WFC decreases sound speed estimation error, and using the estimates for aberration correction improves image resolution and contrast. These promising results have potential to improve pulse-echo imaging for challenging clinical scenarios.
@@ I Introduction
Aberration occurs when an ultrasound beam is not properly focused during beamforming, typically due to heterogeneous sound speed in the media that does not align with the beamforming sound speed. During the image reconstruction process, this impaired focus can cause degradation in the image quality, including decreased resolution, visibility, and contrast [1]. Clinically, aberration can impact medical diagnosis, especially in cases of high depth in vivo imaging, such as abdominal applications [12, 33], and in overweight or obese individuals. To alleviate these problems, aberration correction methods that compensate for the errors due to sound speed heterogeneity have been of high interest for many decades.
@@
Recent approaches to performing distributed aberration correction involve the so-called ”sound-speed-aware” beamformer, where a local sound speed estimate of the imaged medium is obtained, followed by beamforming with time delays calculated using the sound speed estimate to better focus the image throughout the medium [14]. Early pulse-echo local sound speed estimators that could be used for typical medical ultrasound applications were based on solving tomographic inversion problems [15, 28, 34, 26] or estimating the depthwise global (cumulative) sound speed and performing Dix inversion to compute the local sound speed per depth [17, 7]. Although these early methods produced novel estimation results for the traditionally challenging pulse-echo imaging configuration, the estimates often had limited resolution and accuracy, especially when performing in vivo imaging. Nevertheless, from these estimates of the local sound speed, image reconstruction can be used to improve the image focus, often through delay-and-sum (DAS) beamforming but with assumed straight-ray or bent-ray (refracted) propagation paths to compute time delays [3, 9]. Although these methods could be memory- or time-consuming when computing propagation paths at high resolutions, even with not completely accurate local sound speed maps [2, 6], the image quality could have a visible improvement.
@@
To improve upon sound speed inversion methods, iterative closed-loop versions of sound speed estimators have been proposed. An iterative version of pulse-echo tomographic sound speed estimation (iterative CUTE) was developed in [38]. The final predictions from this approach have increased accuracy compared to the prior CUTE model [34] because repeatedly updating the phase shifts for the inversion process using the previously estimated sound speed map alleviates the estimation bias from the initial assumed sound speed. Similarly, [6] improves upon [2], in part by iteratively calculating aberration delays from beamformed data and updating the sound speed estimate in a closed loop. While many of these estimators utilize the relationship between phase shifts, or delays, caused by aberration and the local sound speed in the medium to solve an inversion problem, some of the latest aberration correction approaches optimize an image focus criterion with respect to the local sound speed instead. In earlier decades, optimizing a focus criterion was used for simpler phase-screen models of aberration [19, 21], but has recently been expanded to more distributed correction methods [37]. However, these approaches did not incorporate local sound speed to the image reconstruction process. Recently, we introduced ultrasound autofocusing, a distributed aberration correction technique that ties image focusing criteria to gradient-descent-based sound speed estimation to iteratively update a local sound speed estimate and apply DAS beamforming via delays calculated by a straight-ray model of wave propagation [32]. This method has been shown to produce excellent distributed aberration correction. However, while the sound speed estimation it produces is improved significantly over prior models, the estimation is not entirely accurate because the method does not explicitly optimize for the correct local sound speed, but rather a local sound speed that is optimal for the focusing criterion. Additionally, this method can have difficulty accurately estimating sound speeds for regions with large sound speed differences from the mean global sound speed.
@@
Estimation accuracy can also degrade with data corrupted by acoustic noise. As shown in [17], diffuse reverberation can significantly increase the bias in sound speed estimation due to signal multi-pathing and increased clutter that decreases signal coherence. Additionally, many sound speed estimators utilize a straight- or bent-ray DAS beamforming model to calculate time delays, which may not completely model actual wave propagation through complex heterogeneous media or media with large sound speed changes. However, an alternative beamforming strategy, referred hereafter as wavefield correlation (WFC), can be used to model more complex wave propagation, such as diffraction. Wavefield correlation, also known as shot-profile and reverse-time migration in geophysics [29], consists of simulating the forward propagation of the transmit wavefields and backpropagation of the receive wavefields, followed by the correlation of those wavefields at each propagation step to form the final image axially [30, 8, 2]. This approach can more accurately represent wave propagation during the image reconstruction process and correct for aberration given an estimated local sound speed map [2]. Due to the properties of the angular spectrum, this method can also reduce the effect of diffuse reverberation in the final beamformed data [39] by acting as a spatial matched filter.
@@
In this manuscript, we extend our prior work in [39] to incorporate an image-quality optimization sound speed estimation approach that uses wavefield correlation as the beamformer and the differentiable beamformer approach to improve local sound speed estimates in cluttered and complex media. While this approach is similar to [4] in its usage of WFC as a beamformer and iteratively updating the sound speed model, our work differs fundamentally in the optimization criterion and how the sound speed is updated. We cover the background theory and implementation in Section II and describe simulation, phantom, and in vivo data experiments in Section III that demonstrate improved estimation accuracy and image quality improvements in Section IV.
@@ II Theory
In this section, we explain the wavefield correlation beamforming approach, and then describe how it is integrated into a sound speed optimization method to perform aberration correction and local sound speed estimation. The final process used for estimating the local sound speed is illustrated in Figure 1.
Figure 1: Diagram of the sound speed optimization approach used in this work. The common midpoint phase error is used as the image focus criterion minimized in the optimization objective function. The beamforming process to obtain the focus criterion, using WFC, is tracked using auto-differentiation software, which is then used to backpropagate the loss gradient to update the local sound speed estimation. This iterative gradient-descent-based optimization process then repeats until the estimations have reasonably converged to a local minimum.
@@ II-A Wavefield Correlation Beamformer
Image reconstruction by correlating transmit and receive wavefields has previously been presented for ultrasound applications in [30, 25, 8]. The implementation used in this work follows [2], where transmit pulses are forward propagated and the received data are backward propagated through heterogeneous media using a Fourier domain simulator, and the two wavefields are correlated and summed to form each depth-wise row in the image. We briefly review this approach below, but the reader is referred to Schwab et al. [30] and Ali et al. [8, 2] for a more detailed discussion.
@@ II-A1 Wavefield Simulations
To obtain the transmit and receive wavefields, $p_{tx}(x,z,t)$ and $p_{rx}(x,z,t)$, respectively, for azimuth dimension $x$, depth dimension $z$, and time dimension $t$, the known transmit pulse sequence $p_{tx}(x,z=0,t)$ or channel data $p_{rx}(x,z=0,t)$ at the surface of the transducer is propagated via simulation. Transmit pulses are forward propagated through the heterogeneous media, while the received radiofrequency (RF) data is backward propagated through the same medium. The particular simulator used for this work is a Fourier-domain split-step angular spectrum method (ASM), which is less computationally demanding than a full finite-difference-time-domain (FDTD) simulator but can still model wave propagation through a heterogeneous medium, albeit without multiple reflections [18]. For simplicity, the type of angular spectrum method used in this work does not incorporate attenuation or non-linearity.
To propagate a wavefield $p(x,z,t)$ using the split-step angular spectrum method for a linear transducer, the wavefield is first transformed into the angular spectrum domain through a 2D Fourier transform:
@@
$p_{diffr.}(x,z+\Delta z,f)=\mathcal{F}^{-1}_{k_{x}\longrightarrow x}[p_{diffr.}(k_{x},z+\Delta z,f)]$ (3)
@@
$\begin{split}p(x,z+\Delta z,f)=&exp(\mp j2\pi f\Delta z\Delta u(x,z))\\ &*p_{diffr.}(x,z+\Delta z,f)\,.\end{split}$ (4)
These propagation steps are iterated through successive depths to obtain wavefields for the entire region of interest.
@@ II-A2 Image Reconstruction
$I(x,z)=\int_{-\infty}^{\infty}p_{tx}(x,z,f)p_{rx}^{*}(x,z,f)df\,.$ (5)
This process is equivalent to summing the correlations over all time [8].
II-A3 Beamformer Advantages
This beamforming approach has several advantages that can benefit sound speed estimation and resulting aberration correction. The beamforming process results in a spatio-temporal matched filtering effect from correlating the ideal transmit wavefield with the backpropagated receive data, which can potentially reduce diffuse reverberation noise in the resulting image [39]. Additionally, since a local sound speed map can be used for the simulations, the wave propagation with refractive and diffractive effects of heterogeneous media is more accurately modeled compared to other beamforming approaches, such as DAS using straight ray paths to calculate time delays. Finally, all steps taken during the beamforming process are differentiable, allowing integration into the autofocusing approach we have previously developed for simultaneous aberration correction and sound speed estimation [32].
@@ II-B Gradient-Descent-Based Sound Speed Estimation
Sound speed estimation for this work follows a gradient descent optimization approach proposed in [32]. This method optimizes an image focusing metric relative to the local sound speed map, and updates the sound speed map in the same way as neural network loss backpropagation to minimize the loss function. All operations to calculate the final loss metric involving the local sound speed are tracked in a software package with automatic differentiation functionality, and the computed gradient value from these operations is used to update the local sound speed map. The process then repeats iteratively, much like neural network training, using the updated sound speed map to compute the loss metric and then auto-differentiation to obtain the gradient for updating the sound speed map.
The optimized loss metric used for this work is the combination of the common midpoint phase and the total variation of the sound speed map:
$\mathcal{L}(c)=\mathcal{L}_{CMPE}(c)+\gamma\mathcal{R}_{TV}(c),$ (6)
where $c$ represents the local sound speed map, $\mathcal{L}_{CMPE}$ represents a common midpoint phase error image focus metric, $\gamma$ represents a scalar regularization weight, and $\mathcal{R}_{TV}$ represents anisotropic total variation regularization.
@@ II-B1 Common Midpoint Phase Error
The correlation of beamformed pressure signals between two receive locations, given as
@@
for pressure field $P$ at spatial locations $\boldsymbol{x_{1}}$ and $\boldsymbol{x_{2}}$, is formulated for pulse-echo ultrasound using the van Cittert-Zernike theorem through an aberrating layer in [20] as:
$\begin{split}R(\boldsymbol{x_{1}},\boldsymbol{x_{2}})=&A\cdot e^{j[\phi(\boldsymbol{x_{1}})-\phi(\boldsymbol{x_{2}})]}\cdot e^{j\frac{\pi}{\lambda z}(\boldsymbol{x_{1}x_{1}}-\boldsymbol{x_{2}x_{2}})}\\ &\cdot\int|P_{i}(\boldsymbol{u})|^{2}e^{j2\pi\frac{\boldsymbol{x_{1}}-\boldsymbol{x_{2}}}{\lambda z}\boldsymbol{u}}\mathop{}\!\mathrm{d}u.\end{split}$ (8)
In this equation, $A$ is the resulting amplitude term, $\phi$ is the aberration expressed as a phase, $\lambda$ is the wavelength, $z$ is the depth, and $P_{i}$ is the transmitted pressure field. The integral term represents the Fourier transform of the source pressure field intensity, per the van Cittert-Zernike theorem. The second exponential term outside of the Fourier transform represents a location-dependent phase term based on the Fresnel approximation. The first exponential then represents any remaining phase difference between the two receive locations due to aberration. When the two receive locations share a common midpoint about the transducer axis, i.e.,
@@
and the data is beamformed, all phase terms except the first term cancel out or become deterministic. The first phase term thus accounts for any beamforming inaccuracies that cause aberration in the final beamformed data. Under perfect focusing, the first phase term should reduce to 0. The mean phase difference between common midpoint signals, called the common midpoint phase error [32], is proposed as a measure of the focusing quality of the beamformer and utilized as an optimization objective, i.e.,
$\mathcal{L}_{CMPE}=\angle R(\boldsymbol{x_{1}},\boldsymbol{x_{2}}).$ (11)
In Simson et al. [32], the common midpoint phase error for the beamformed data over all common midpoint subapertures of size 17 elements was used after excluding values where the correlation coefficient at that location was lower than 0.9 to exclude noisy signals. Similarly, the work in this paper uses this filtered version of the common midpoint phase error (CMPE), which is minimized during the optimization process. This metric is computed through software beamforming of the data using WFC. Separate simulations are performed for each subaperture considered, and both the ASM simulations and correlation process for final image formation are composed of differentiable operations.

### v2

@@ Abstract
In pulse-echo ultrasound, aberration often degrades image quality when beamforming does not account for wavefront distortions. To address this issue, local sound speed estimators have been developed in the past decade for distributed aberration correction. Recently, methods based on iterative optimization have improved sound speed accuracy with respect to earlier approaches. However, the accuracy of these newer methods is limited by media with reverberation clutter and by the straight-ray model of wave propagation. To address these challenges, we propose using wavefield correlation (WFC) beamforming when performing sound speed optimization. WFC, an ultrasound adaptation of reverse time migration, correlates simulated forward-propagated transmit wavefields and backwards-propagated receive wavefields in order to reconstruct images. This process more accurately models wave propagation in heterogeneous media and can decrease diffuse clutter due to its spatiotemporal matched filtering effect. We implement herein a WFC beamformer using an auto-differentiation software and estimate the sound speed map by optimizing a regularized common-midpoint phase focusing criterion using gradient descent. This approach is compared to a previous method relying on delay and sum (DAS) with straight-ray time delay calculations on a variety of simulated, phantom, and in vivo data with large sound speed variations and clutter. Results show that using WFC decreases sound speed estimation error, leading to improvements in resolution and contrast in the corrected image. In particular, these promising results have potential to improve pulse-echo imaging for challenging clinical scenarios.
@@ I Introduction
In ultrasound B-mode imaging, aberrations occur when a beam is not optimally focused during beamforming, typically due to spatial variations in sound speed within the medium that differ from the constant sound speed assumed for image reconstruction. This impaired focus can cause degradation in the image quality, including decreased resolution, visibility, and contrast [1]. Clinically, aberration can impact medical diagnosis, especially when imaging at high depths, such as in abdominal screening [13, 36] and with overweight or obese individuals. To alleviate these problems, aberration correction methods that compensate for the focusing errors caused by sound speed heterogeneity have been of high interest for many decades.
@@
Recent state-of-the-art approaches for distributed aberration correction involve the so-called ”sound-speed-aware” beamformer, where a local sound speed estimate of the imaged medium is obtained, followed by beamforming with time delays calculated using the sound speed estimate to better focus the image throughout the medium [15]. Early pulse-echo local sound speed estimators that could be used for typical medical ultrasound applications were based on solving tomographic inverse problems [16, 31, 37, 29] or estimating the depthwise global (cumulative) sound speed and performing Dix inversion to compute the local sound speed per depth [18, 7]. These early methods produced novel sound speed estimation results for the traditionally challenging pulse-echo imaging configuration and greatly advanced knowledge of local sound speed estimation in pulse-echo ultrasound, but the estimates often were limited in resolution and accuracy, especially in vivo. Nevertheless, from these estimates of the local sound speed, image reconstruction can be performed to improve the image focus, often through delay-and-sum (DAS) beamforming but with assumed straight-ray or bent-ray (refracted) propagation paths to compute time delays [41, 3, 9]. Although these methods could be memory- or time-consuming when computing propagation paths at high resolutions, they often lead to significant improvements in visual quality, even with inaccuracies in the computed local sound speed maps [2, 6].
@@
To improve the determination of local sound speed, iterative closed-loop versions of sound speed estimators have been proposed. An iterative version of pulse-echo tomographic sound speed estimation (iterative CUTE) was developed in [42]. The final predictions from this approach have increased accuracy compared to the prior CUTE model [37] because repeatedly updating the phase shifts for the inversion process using the previously estimated sound speed map alleviates the estimation bias from the initial assumed sound speed. Likewise, [6] improves upon [2], in part by iteratively calculating aberration delays from beamformed data and updating the sound speed estimate in a closed loop. While many of these estimators utilize the relationship between measured phase shifts, or aberration delays, caused by local sound speed variations in the medium to solve an inverse problem, some of the more recent aberration correction approaches optimize an image focusing criterion with respect to the local sound speed instead. In earlier decades, optimizing a focusing criterion was used for simpler phase-screen models of aberration [21, 24], but this principle has recently been expanded to target distributed aberration correction as well [40]. However, these approaches do not incorporate local sound speed to the image reconstruction process. Recently, we introduced ultrasound autofocusing, a distributed aberration correction technique that ties image focusing criteria to gradient-descent-based sound speed estimation to iteratively update a local sound speed estimate and apply DAS beamforming via delays calculated by a straight-ray model of wave propagation [35]. In particular, this method has been shown to produce excellent distributed aberration correction.
@@
While sound speed estimation improved significantly with respect to prior models, the accuracy of ultrasound autofocusing for local sound speed imaging is still limited by several factors. First, estimation accuracy can degrade with data corrupted by acoustic noise. As shown in [18], diffuse reverberation can significantly increase the bias in sound speed estimation due to multiple scattering and increased clutter that decreases signal coherence. Additionally, many sound speed estimators utilize a straight- or bent-ray DAS beamforming model to calculate time delays which do not factor in diffraction effects, leading to inaccuracies in complex heterogeneous media or media with large sound speed changes. However, an alternative beamforming strategy, referred hereafter as wavefield correlation (WFC), can be used to model more complex wave propagation such as diffraction. Wavefield correlation, also known as shot-profile and reverse-time migration in geophysics [32], consists of simulating the forward propagation of the transmit wavefields and backpropagation of the receive wavefields, followed by the correlation of those wavefields to form the final image [33, 8, 2]. This approach can more accurately represent wave propagation during the image reconstruction process and correct for aberration given an estimated local sound speed map [2]. Due to the properties of the wavefield simulation and subsequent correlation, this method can also reduce the effect of diffuse reverberation in the final beamformed data [43] by acting as a spatial matched filter.
@@
In this manuscript, we extend our prior work in [43] to incorporate sound speed estimation based on image-quality optimization using WFC as the beamformer in the autofocusing approach [35], with the goal of improving local sound speed estimation in complex media. While this approach is similar to [4] in its usage of WFC as a beamformer and iteratively updating the sound speed model, our work differs fundamentally in the optimization criterion and how the sound speed estimate is updated. We review the background theory and the adaptation of WFC to the autofocusing model and then demonstrate the application of this method to simulation, phantom, and in vivo imaging.
@@ II Theory
In this section, we explain the wavefield correlation beamforming approach, and describe how it is integrated into a sound speed optimization method to perform aberration correction and local sound speed estimation. The overall process used for estimating the local sound speed is illustrated in Figure 1.
Figure 1: Diagram of the sound speed optimization approach. The common midpoint phase error is used as the image focusing criterion that we seek to optimize. The WFC beamforming process to obtain the focus criterion is tracked using auto-differentiation software, which is then used to backpropagate the loss gradient to update the local sound speed estimation. This iterative gradient-descent-based optimization process loops until a convergence criterion is met.
@@ II-A Wavefield Correlation Beamformer
Image reconstruction by correlating transmit and receive wavefields has previously been presented for ultrasound applications in [33, 28, 8]. The implementation used in this work follows [2], where transmit pulses are forward propagated and the received data are backward propagated through heterogeneous media using a Fourier split-step simulator, and the two wavefields are correlated and summed to form each constant-depth row in the image. We briefly review this approach below, but we refer the reader to Schwab et al. [33], Ali [8], and Ali et al. [2] for a more detailed discussion.
@@ II-A1 Wavefield Simulations
In the Wavefield Correlation (WFC) beamformer, the transmit ($p_{\text{tx}}(x,z,t)$) and receive ($p_{\text{rx}}(x,z,t)$) wavefields for azimuth dimension $x$, depth dimension $z$, and time dimension $t$ are required. These wavefields are obtained from the known or approximated transmit pulse $p_{\text{tx}}(x,z=0,t)$ and received channel data $p_{\text{rx}}(x,z=0,t)$ of the transducer, which are propagated via simulation. Transmit pulses are forward propagated through the heterogeneous media, while the received radiofrequency (RF) channel data are backward propagated through the same medium. The simulator used in this approach is a Fourier-domain split-step angular spectrum method (ASM), which is less computationally demanding than a full finite-difference-time-domain (FDTD) simulator typically used in reverse time migration. ASM can model wave propagation through a heterogeneous medium, albeit without considering multiple reflections [19]. For simplicity, the implementation of ASM used herein does not incorporate attenuation or non-linearity.
Let us restrict ourselves to linear transducers. To propagate a wavefield $p(x,z,t)$ using ASM, the wavefield is first transformed into the angular spectrum domain through a 2D Fourier transform:
@@
$p_{\text{diffr.}}(x,z+\Delta z,f)=\mathcal{F}^{-1}_{k_{x}\longrightarrow x}[p_{\text{diffr.}}(k_{x},z+\Delta z,f)]$ (3)
@@
$\begin{split}p(x,z+\Delta z,f)=&\exp(\mp j2\pi f\Delta z\Delta u(x,z))\\ &*p_{\text{diffr.}}(x,z+\Delta z,f)\,.\end{split}$ (4)
These propagation steps are iterated through successive depths to obtain wavefields for the entire insonified medium.
@@ II-A2 Image Reconstruction
$I(x,z)=\int_{-\infty}^{\infty}p_{\text{tx}}(x,z,f)p_{\text{rx}}^{*}(x,z,f)df\,.$ (5)
This process is equivalent to integrating the correlations over all time instances [8].
This beamforming approach has several advantages that benefits sound speed estimation and the resulting aberration correction. Through correlation of an ideal transmit wavefield with the backpropagated receive data, this beamforming process results in a spatio-temporal matched filtering effect, which can reduce diffuse reverberation noise in the resulting image [43]. Additionally, since a local sound speed is inherently assumed during wavefield simulation, wave propagation with refractive and diffractive effects in heterogeneous media is more accurately modeled compared to other beamforming approaches, such as DAS using straight ray paths to calculate time delays. Finally, all steps taken during the beamforming process are differentiable, allowing integration into the autofocusing approach we have previously developed for simultaneous aberration correction and sound speed estimation [35].
@@ II-B Gradient-Descent-Based Sound Speed Estimation
In this work, we rely on the gradient descent optimization approach proposed in [35] for sound speed estimation. This method optimizes an image focusing metric relative to the local sound speed map, and updates the sound speed map in the same way as neural network loss backpropagation to minimize the loss function. All operations to calculate the final loss metric involving the local sound speed are tracked in a software package with automatic differentiation functionality, and the computed gradient value from these operations is used to update the local sound speed map. The process then repeats iteratively, much like neural network training, using the updated sound speed map to compute the loss metric and then auto-differentiation to obtain the gradient for updating the sound speed map.
The optimized loss metric used in this work is the sum of the common midpoint phase and a regularization term:
$\mathcal{L}(c)=\mathcal{L}_{\text{CMPE}}(c)+\gamma\mathcal{R}_{\text{TV}}(c),$ (6)
where $c$ represents the local sound speed map, $\mathcal{L}_{\text{CMPE}}$ represents the common midpoint phase error image focus metric. Here, we follow [35] and use an anisotropic total variation regularizer $\mathcal{R}_{\text{TV}}$ with strength parameterized by a positive scalar regularization weight $\gamma$.
@@ II-B1 Common Midpoint Phase Error
The correlation of beamformed pressure signals at two receive locations $\boldsymbol{x_{1}}$, $\boldsymbol{x_{2}}$ is given by
@@
where $P$ denote the pressure signal and $\langle\cdot\rangle$ denotes cross-correlation. Using the van Cittert-Zernike theorem, the correlation $R$ is formulated for pulse-echo ultrasound through an aberrating layer according to [22, 23] as:
$\begin{split}R(\boldsymbol{x_{1}},\boldsymbol{x_{2}})=&A\cdot e^{j[\phi(\boldsymbol{x_{1}})-\phi(\boldsymbol{x_{2}})]}\cdot e^{j\frac{\pi}{\lambda z}(\boldsymbol{x_{1}x_{1}}-\boldsymbol{x_{2}x_{2}})}\\ &\cdot\int_{U}|P_{i}(\boldsymbol{u})|^{2}e^{j2\pi\frac{\boldsymbol{x_{1}}-\boldsymbol{x_{2}}}{\lambda z}\boldsymbol{u}}\mathop{}\!\mathrm{d}\boldsymbol{u}.\end{split}$ (8)
In (8), $A$ is the resulting amplitude term, $\phi$ is the aberration expressed as a phase, $\lambda$ is the wavelength, $z$ is the depth, and $P_{i}$ is the transmitted pressure field, which insonifies the medium $U$. The integral term represents the Fourier transform of the source pressure field intensity, per the van Cittert-Zernike theorem. The second exponential term outside of the Fourier transform represents a location-dependent phase term based on the Fresnel approximation. The first exponential then represents any remaining phase difference between the two receive locations due to aberration. When the two receive locations share a common midpoint about the transducer axis, i.e.,
@@
and the data are beamformed, all phase terms except the first term cancel out or become deterministic. The first phase term thus accounts for any beamforming inaccuracies that cause aberration in the final beamformed data. Under perfect focusing, the first phase term should reduce to 0. The mean phase difference between common midpoint signals, called the common midpoint phase error [35], is proposed as a measure of the focusing quality of the beamformer and utilized as an optimization objective, i.e.,
$\mathcal{L}_{\text{CMPE}}=\angle R(\boldsymbol{x_{1}},\boldsymbol{x_{2}}).$ (11)
In Simson et al. [35], the common midpoint phase error for the beamformed data over all common midpoint subapertures of size 17 elements was used after excluding values where the correlation coefficient at that location was below a threshold $\zeta$ to exclude noisy signals (e.g. from anechoic regions). Similarly, the present work uses this filtered version of the common midpoint phase error (CMPE), which is minimized during the optimization process. This metric is computed through software beamforming of the data using WFC. Separate simulations are performed for each subaperture considered, and both the ASM simulations and correlation process for final image formation are composed of differentiable operations.